# Co-VisiON: Co-Visibility ReasONing on Sparse Image Sets of Indoor Scenes

### v1

@@ 4.1 Naive Baselines
We first benchmark several conventional approaches for Co-VisiON, where inference involves exhaustive pairwise comparisons between all node pairs to generate a topology map. This includes binary classification, supervised contrastive learning and feature matching.
Pairwise supervised classification. We test three classic image classification backbones, including ViT [15], VGG [50], and ResNet [21]. They extract representations for a pair of images and append a classification head to predict binary co-visibility probabilities. The models are fine-tuned end-to-end. For inference, we threshold the pairwise prediction probabilities to generate a binary adjacency matrix for each scenario. The area under the curve (AUC) is calculated at various thresholds ranging from 0 to 1, as detailed in Sec. 3.2.
Contrastive learning approaches. We test two contrastive approaches. The first adopts a ResNet backbone with the InfoNCE loss [38] following the architecture of SimCLR [10]. Unlike SimCLR, we conduct supervised contrastive learning. Pairs of images connected in the co-visibility graph are treated as positive pairs, and others as negative pairs. The model is fine-tuned end-to-end from the pretrained SimCLR. The second approach utilizes NetVLAD [4], a widely used method in visual place recognition. It uses a ResNet backbone followed by a VLAD aggregation layer to extract global image descriptors. For both approaches, the pairwise feature distance, normalized by the maximum feature distance (ranging from 0 to 1), is used during inference to predict the edge probability for the co-visibility graph.
Image feature matching. SIFT+RANSAC [18] and SuperGlue [44] are used as baseline methods for image feature matching due to their classic and widespread use. SuperGlue processes image pairs to detect key points and compute matches via a graph neural network, while SIFT+RANSAC employs traditional feature matching and outlier filtering. The percentage of matches compared to the maximum possible matches is treated as the edge probability for the co-visibility graph.
Discussion. We present the experiment results in  Tab. 2. Compared with only reporting the best IoU, we also emphasize the importance of IoU invariance across various thresholds as a critical performance metric, particularly in real-world applications where real-time threshold optimization may not be feasible. A robust baseline should demonstrate consistent performance across a wide range of thresholds, ensuring acceptable results regardless of the threshold chosen. In Fig. 4, we plot the AUC curves of all baselines. Vision Transformer (ViT) stands out for its stable IoUs across diverse thresholds among all the conventional baselines. Contrastive learning shows a similar trend. Comparing the baselines by their categories, supervised learning, particularly with ViT, consistently outperforms feature-matching methods in both datasets shown in Tab. 2. Contrastive learning methods, although trailing behind supervised methods, still outperform feature-matching techniques by effectively managing positive and negative pairs in feature space. This showcases the great potential of contrastive learning in addressing co-visibility issues. All conventional baselines are substantially worse than the human evaluation. Pairwise image matching and classification baselines are not sufficient to solve Co-VisiON.
@@ 4.2 Improved Baselines
Going beyond conventional baselines, we explore whether recent vision models can perform better on Co-VisiON, including state-of-the-art 3D stereo reconstruction methods that can inherently derive co-visibility, and large Vision and Language Models (VLMs) that excel in visual understanding. More importantly, we propose and test a new baseline–Covis, leveraging multi-view masked classification.
VLM for Co-VisiON. We leverage external API calls to harness the reasoning capabilities of state-of-the-art large vision and language models (VLMs). This method encodes images into base64 format and sends them to GPT-4o [39] and Qwen2.5-VL 72b [64] models. Then they use their pre-trained capabilities to compare the images, compute pairwise scores, and generate a topology map. We further include results from recent Spatial VLM methods like SpatialRGPT [11] and Gemini-2.0-Flash [57] in the appendix.
3D Reconstruction for Co-VisiON. We use the DUSt3R and MV-DUSt3R+ frameworks to generate global point clouds from input photos. The presence of binary topology edges is determined by the overlap of these point clouds. This approach is counterintuitive, as 3D reconstruction typically relies on topology map node relationships. Instead, we use 3D reconstruction to generate the topology map to highlight its limitations, including domain gap issues and frequent misjudgments of negative samples shown in Tab. 2.
@@ 4.3 Covis
Architecture. Both pairwise and multi-view Covis use CroCov2 [60] as the backbone with a BCE loss to predict the relationship between images. In the pairwise setting, a reference image is paired with a single image from another viewpoint, and the BCE loss predicts whether the two images are co-visible. In contrast, the multi-view Covis is built on the MV-DUSt3R [55] encoder-decoder backbone that facilitates the exchange of information between multiple images from diverse viewpoints, allowing for richer spatial reasoning. During training, positive and negative pairs are constructed based on the ground truth co-visibility graph, and during inference, a mixture of positive and negative pairs from the same scene is used to assist pairwise reasoning. The multi-view approach captures broader context by processing multiple images simultaneously.
Mask prediction. We found that non-covisible information contaminates the learned representation of the reference image, hindering effective learning from positive views (see Appendix for details). To address this, we introduce a learnable mask $M\in\mathbb{R}^{H\times W}$, jointly optimized as an auxiliary objective and used to filter out irrelevant features.
$\mathbf{M}=\operatorname{conv}(\mathbf{F}),$ (3)
where $F\in\mathbb{R}^{H\times W\times C}$ denotes the pixel-wise feature map obtained by projecting the decoder output. After obtaining $\mathbf{M}$, we apply it to $\mathbf{F}$ via a dot product at each pixel, preserving relevant information from the positives:
$\mathbf{F}^{\prime}=\mathbf{F}\odot\mathbf{M},$ (4)
where $\odot$ denotes pixel-wise multiplication. Both the topology generation and mask prediction tasks use BCE loss but serve distinct purposes: the former predicts spatial relationships based on flattened 1D global features for image pairs, while the latter supervises pixel-wise co-visible regions, providing stronger localized signals.
@@
4.4 Discussion
@@
Category Method Backbone Pairwise(✗) / Multiview(✓) Gibson HM3D
@@
VLM GPT-4o [39] 0.63 0.63 0.59 0.59
@@
Pairwise vs multi-view performance. Across both 3D reconstruction and Covis with mask prediction, we observe that multi-view approaches consistently outperform pairwise approaches by about 3% in IoU, as shown in  Tab. 2. Multi-view inputs provide richer spatial context, leading to more accurate 3D reconstruction and better feature learning. In contrast, pairwise Covis is limited by the information in a single image pair, resulting in reduced performance.
The importance of masking. As discussed, multi-view methods typically outperform pairwise methods. However, our experiments with the non-masking version of Covis reveal that their performance degrades sharply as the proportion of negatives increases, particularly in multi-view settings. This is due to feature contamination, where irrelevant features from positives and negatives interfere during information exchange. Inspired by the confidence mechanism in DUSt3R [58], we introduce a learnable mask to filter out irrelevant information. Our results show that applying this mask improves performance by approximately $10\%$, with especially pronounced gains in multi-view scenarios.
VLM performance. We test both proprietary and open-source models without fine-tuning. GPT-4o achieves 63% IoU, which is currently the closest to the human reasoning performance of 72%, top-performed in  Tab. 2. This suggests that reasoning-capable models can better identify co-visibility, especially under the sparse conditions of the Co-VisiON benchmark. In contrast, the open-source Qwen2.5-VL 72b performs less favorably, reaching only 41%, even lower than many visual baselines. This highlights the instability of VLMs on unseen tasks, despite their promising potential for vision understanding.
@@ 5.1 DUSt3R for scene reconstruction
Table 3: Comparison of different graph structures within the DUSt3R framework. We evaluate the Average Pose Error (APE), along with time complexity (in seconds) and memory cost (in GB) per scene. Results are averaged over the Gibson test set. The Random graph fails in the 3D reconstruction test, making APE not applicable.
@@
Quantitative Analysis of Different Graphs. We calculate the average pose errors (APE) by matching estimated poses with their closest ground truth poses to evaluate pose estimation and reconstruction quality, as shown in Tab. 3. As expected, using complete and ground truth (GT) graphs yields lower APE, demonstrating more accurate 3D reconstruction. However, obtaining a ground truth graph and using the complete graph incurs high annotation and computational costs. Moreover, random and star graphs fail to produce a reliable 3D map, as shown in Fig. III. Therefore, we find the co-visibility graph to be one of the most suitable options for DUSt3R, balancing memory costs with a reliable 3D map, as shown in Fig. 5. Detailed reconstruction visualizations are available in the supplementary.
@@ 5.2 Automated Training Set Labeling
Figure 6: Visualizations of the reconstructed co-visibility regions during inference. The top row shows model inference using the co-visibility graph for training, while the bottom row uses the high co-visibility graph. The three columns represent the reference image, predicted image, and ground truth target image. In the predicted images, the masked regions are visible as network input, and the brightened areas indicate the reconstructed output. Additionally, the random graph fails in cross-view completion.
@@ 6 Conclusion
Limitation. The current co-visibility reasoning task presents several notable limitations. For the dataset, the scenes are restricted to the Gibson and HM3D environment, and the human annotations are limited. With the proposed data collection pipeline and website, our future work will continue to enlarge and diversify the dataset scenes with other sources, including the real world.
Summary. This paper proposes the Co-Visibility reasONing (Co-VisiON) task, where models are presented with sparse image sets taken in an indoor environment and are tasked with reasoning their co-visibility relationships by determining view connectivity and predicting a binary co-visibility graph. We curate a dataset designed for this task, containing both human and automated annotations, and conduct comprehensive experiments to show how the Co-VisiON task challenges dominant image classification and matching baselines. Additionally, we introduce Covis, which is able to leverage multi-view reasoning to effectively capture complex spatial relationships between images, outperforming traditional pairwise approaches. We hope that the Co-VisiON task can facilitate future research in embodied scene understanding.
Acknowledgement. We thank Kevin Lee, Kingston Yip, and Mandy Mao for their early investigation on the dataset and baselines, and Leon Zhang and Ammaar Firozi for their contributions to the dataset. The work was supported in part through NSF grants 2238968, 2322242, 2121391, and 2036870, and the NYU IT High Performance Computing resources, services, and staff expertise.
@@
This supplementary provides additional details and results that could not fit in the main paper. Specifically, we include: (1) dataset and annotation generation (Gibson and HM3D), (2) top-down visualization of co-visibility graph, (3) evaluation metric details, (4) graph definitions used in DUSt3R and CroCo experiments, (5) ablation studies on Covis, (6) additional models and failure mode analysis, and (7) Sim2Real downstream task details. All experiments and dataset generation were conducted on an HPC cluster with A100 and V100 GPUs, using a single GPU per run.
@@ B.1.2 Human co-visibility annotation
We will release both the automatically labeled and the manually labeled datasets to facilitate future research. The human annotation website will also be released so that anyone can extend the co-visibility reasoning task to a larger scale and for more diverse scenarios.
@@ G Ablation Study of CoVis
G.1 Effect of noisy mask on Covis
Mask prediction plays a crucial role in CoVis, as demonstrated in  Tab. 2. In this subsection, we explore how the accuracy of ground truth co-visible masks—ranging from accurate to noisy—affects the model’s performance on the test set.
To simulate noisy supervision, we introduce boundary-localized Gaussian perturbations to the ground truth binary masks. First, a Laplacian filter is used to detect boundaries between foreground and background. Gaussian noise is then added to these boundary pixels, with standard deviations of $\sigma=0.01$ for low noise and $\sigma=0.5$ for high noise. The perturbed mask is re-binarized using a threshold of $\tau=0.5$, while retaining all original positive pixels to avoid removing valid co-visible regions. If the original mask is entirely empty, no noise is applied.
When provided with low levels of noise in the co-visibility masks, the performance on both IoU and AUC remains comparable or even slightly improves due to the regularizing effect of noise, akin to data augmentation. However, as shown in Tab. I, excessive noise leads to a significant drop in performance, highlighting the importance of maintaining reasonable mask quality.
Table I: Comparison of different mask types with varying noise levels in Covis’ test performance (IoU means Graph IoU).
Mask Type Noise Gibson HM3D
IoU AUC IoU AUC
No Mask - 0.52 0.49 0.48 0.42
Supervised Mask None 0.59 0.57 0.57 0.56
Supervised Mask Low 0.59 0.56 0.59 0.55
Supervised Mask High 0.56 0.50 0.50 0.48
Applied GT Mask - 0.74 0.72 0.71 0.69
G.2 Zero-shot Evaluation
We further evaluate the generalization ability of our Covis by training on one dataset (either Gibson or HM3D) and testing on both Gibson and HM3D individually. As shown in Tab. II, Multi-view Covis consistently outperforms pairwise Covis in both in-domain and zero-shot settings, demonstrating better robustness to dataset’s domain shifts. We also evaluate the predicted masks using the standard IoU metric (not graph IoU), and achieve an average IoU of 67.3% against the binarized ground-truth masks.
Table II: Zero-shot evaluation of multi-view Covis and pairwise Covis across different training and test sets.

### v3

@@ 4.1 Naive Baselines
We first benchmark several conventional approaches for Co-VisiON, where inference involves exhaustive pairwise comparisons between all node pairs to generate a co-visibility map. This includes binary classification, supervised contrastive learning and feature matching.
Pairwise supervised classification. We evaluate three classic backbones—ViT [14], VGG [49], and ResNet [20]—for pairwise co-visibility prediction. Each model extracts image pair features and uses a classification head to predict binary probabilities, trained jointly. During inference, predictions are thresholded to form a binary adjacency matrix. AUC is computed across thresholds from 0 to 1 (detailed in Sec. 3.2).
Contrastive learning approaches. We test two contrastive approaches. The first uses a ResNet backbone with InfoNCE loss [37], following SimCLR [10], but with supervised contrastive learning: co-visible image pairs are positives, others are negatives. The model is fine-tuned from pretrained SimCLR. The second uses NetVLAD [5], which combines a ResNet backbone with a VLAD aggregation layer to extract global descriptors. For both methods, pairwise feature distance (normalized to [0, 1]) is used to predict edge probabilities in the co-visibility graph.
Image feature matching. SIFT+RANSAC [17] and SuperGlue [43] serve as classic baselines for image feature matching. SuperGlue uses a graph neural network to match keypoints between image pairs, whereas SIFT+RANSAC relies on traditional feature matching with outlier rejection. The percentage of matches compared to the maximum possible matches is treated as the edge probability for the co-visibility graph.
@@ 4.2 Improved Baselines
Beyond conventional baselines, we explore whether recent vision models can perform better on Co-VisiON, including state-of-the-art 3D stereo reconstruction methods that can inherently derive co-visibility, and large Vision Language Models (VLMs) that excel in visual understanding. More importantly, we propose and test a new baseline–Covis, leveraging multi-view masked classification.
Figure 4: An example of the prompt used in our experiments. A unified prompt is adopted across all vision-language models.
Large VLMs for Co-VisiON. We evaluate the visual reasoning capabilities of general-purpose VLMs, including GPT-4o [38], Gemini 2.0 Flash [56], and Qwen2.5-VL 72b [63]. Each model takes as input an image pair along with a prompt that explicitly asks it to reason about the co-visibility between the two views. In addition, we employ SpatialRGPT [11], a model specifically tailored for spatial reasoning tasks. SpatialRGPT builds on VILA [30], which natively supports multi-image input. We extend it to process image pairs with a shared visual encoder and a fine-tuned LLaMA3-70B [3] decoder for co-visibility reasoning. We use the same prompt for all VLM methods, as shown in Fig. 4. These models leverage their pre-trained reasoning capabilities to compare the images, compute pairwise co-visibility scores, and generate a co-visibility graph.
3D Reconstruction models for Co-VisiON. We use the DUSt3R [57] and MV-DUSt3R+ [54] frameworks to generate global point clouds from input photos. The presence of binary topology edges is determined by the overlap of these point clouds. This approach is counterintuitive, as 3D reconstruction typically relies on topology map node relationships. Instead, we use 3D reconstruction to generate the co-visibility graph to highlight its limitations, including domain gap issues and frequent misjudgments of negative samples shown in Tab. 4.
@@ 4.3 Covis
Architecture. CroCov2 [59] serves as the backbone for both pairwise and multi-view Covis, trained with a BCE loss to predict co-visibility. Experiments show that freezing the backbone during training yields the best results; this variant is referred to as Covis-freeze in subsequent discussions. In the pairwise setting, a reference image is paired with one other viewpoint to predict co-visibility. The multi-view Covis extends this by using the MV-DUSt3R [54] encoder-decoder, enabling information exchange across multiple views for richer spatial reasoning. During training, positive and negative pairs are sampled from the ground truth co-visibility graph, while inference uses mixed pairs from the same scene to aid pairwise prediction. This multi-view design captures broader context by processing multiple images simultaneously.
Mask prediction. Non-co-visible information contaminates the reference image’s representation, hindering positive view learning (see Appendix). To filter out irrelevant features, we introduce a learnable mask $\mathbf{M}^{i}\in\mathbb{R}^{H\times W}$ for each view $i$, computed as $\mathbf{M}^{i}=\operatorname{conv}(\mathbf{X}^{i})$, where $\mathbf{X}^{i}\in\mathbb{R}^{H\times W\times C}$ is a pixel-wise feature map projected from the corresponding decoder token $\mathbf{F}^{i}_{d}\in\mathbb{R}^{T}$ through a shared MLP: $\mathbf{X}^{i}=\operatorname{MLP}(\mathbf{F}^{i}_{d})$. The MLP learns to expand 1D token into a 2D spatial feature map and is shared across all views except the reference view.
We then apply the mask to $\mathbf{X}$ via pixel-wise multiplication to preserve relevant regions from positive views: $\mathbf{X}^{\prime}=\mathbf{X}\odot\mathbf{M}$, where $\mathbf{X}^{\prime}$ denotes the masked feature map, and $\odot$ denotes pixel-wise multiplication. Both the co-visibility prediction and mask prediction tasks use BCE loss but serve distinct purposes: the former predicts spatial relationships based on flattened 1D global features for image pairs, while the latter supervises pixel-wise co-visible regions, providing stronger localized signals.
Table 2: Comparison of different mask types with varying noise levels in Covis’ test performance (IoU means Graph IoU).
Mask Type Noise Gibson HM3D
IoU AUC IoU AUC
No Mask - 0.52 0.49 0.48 0.42
Supervised Mask None 0.59 0.57 0.57 0.56
Supervised Mask Low 0.59 0.56 0.59 0.55
Supervised Mask High 0.56 0.50 0.50 0.48
Applied GT Mask - 0.74 0.72 0.71 0.69
@@
5 Discussion
Overall, our experiments demonstrate that GPT-4o and CoVis consistently outperform both naive and enhanced baselines across diverse datasets. In this section, we first discuss key findings observed throughout the experiments, followed by an analysis of common failure modes to identify potential avenues for improvement.
5.1 Key findings
Summary of Naive baseline performance. We present results in Tab. 4 and AUC curves in Fig. 5. Among conventional baselines, Vision Transformer (ViT) achieves the most stable IoUs across thresholds, with contrastive learning showing a similar trend. Supervised methods, especially ViT, outperform feature-matching approaches on both datasets. Contrastive learning, though trailing supervised methods, still outperforms feature-matching by effectively managing positive and negative pairs—highlighting its potential for co-visibility reasoning. All conventional baselines fall significantly short of human performance. Pairwise matching and classification methods alone are insufficient for Co-VisiON.
Table 3: Best IoU performance across different difficulty levels. The top-3 performers under each difficulty level are color-coded.
Setting Method Easy Med. Hard Avg.
Image Overlap (edge level) GPT-4o 0.97 0.92 0.34 0.63
Gemini-2.0-Flash 1.00 0.99 0.14 0.42
Qwen2.5-VL 72b 1.00 0.99 0.14 0.41
SpatialRGPT 1.00 0.99 0.13 0.35
Covis 0.99 0.88 0.24 0.59
Covis (freeze) 1.00 0.89 0.30 0.61
Scene Sparsity (graph level) GPT-4o 0.83 0.64 0.57 0.63
Gemini-2.0-Flash 0.75 0.39 0.28 0.42
Qwen2.5-VL 72b 0.77 0.40 0.28 0.41
SpatialRGPT 0.72 0.38 0.26 0.35
Covis 0.80 0.63 0.52 0.59
Covis (freeze) 0.81 0.65 0.54 0.61
@@ 5.1 Key findings
Category Method Backbone
Pairwise(✗) /
Multiview(✓)
Gibson HM3D
@@
Covis-freeze (Ours) CroCo v2 [59] 0.61 0.57 0.58 0.56
VLM Qwen2.5-VL 72b [63] 0.41 0.41 0.39 0.39
Gemini-2.0-Flash [56] 0.42 0.42 0.39 0.39
SpatialRGPT [11] 0.49 0.49 0.37 0.37
@@
Pairwise vs multi-view performance. Across both 3D reconstruction and Covis, we observe that multi-view approaches consistently outperform pairwise approaches by about 3% in IoU, as shown in  Tab. 4. Multi-view inputs provide richer spatial context, leading to more accurate 3D reconstruction and better feature learning. In contrast, pairwise Covis is limited by the information in a single image pair, resulting in reduced performance.
The importance of masking. Multi-view methods generally outperform pairwise ones. However, without masking, Covis suffers a sharp performance drop as the negative sample ratio increases, particularly in multi-view scenarios. This is because more negative information is introduced, acting as distractors during training. More specifically, irrelevant positive and negative features exchanged during the encoder-decoder process lead to feature contamination. In Tab. 2, the learnable mask helps filter out irrelevant regions and consistently improves performance, achieving up to a  10% gain in challenging multi-view settings.
To further examine the effect of mask quality, we simulate noisy supervision by injecting boundary-localized Gaussian noise into ground truth masks. Boundaries are detected via Laplacian filtering; noise with standard deviation $\sigma\in{0.01,0.5}$ is added and binarized at threshold $\tau=0.5$, while preserving all original positives. Low-level noise slightly improves performance due to regularization, while high noise degrades it significantly, highlighting the importance of accurate masks.
VLM performance. We evaluate proprietary and open-source models without fine-tuning. GPT-4o attains 63% IoU-closest to human performance at 72% and best in Tab. 4 - demonstrating superior co-visibility reasoning capability under sparse conditions. Conversely, open-source models Qwen2.5-VL 72b and Gemini-2.0-Flash perform worse, achieving just over 40%, below many visual baselines. Moreover, as a spatial reasoning specialist, SpatialRGPT scores under 50%, highlighting VLM instability on novel tasks despite their visual understanding capabilities.
5.2 Failure Mode Analysis
To further assess the robustness of each method, we conduct a failure mode analysis by categorizing the evaluation data along two dimensions: image overlap and scene sparsity. These correspond to different levels of sparsity in the underlying co-visibility graph: image overlap reflects edge-level sparsity between image pairs, while scene sparsity captures graph-level sparsity across the entire scene. This enables a more fine-grained understanding of model behavior when either the overlap between image pairs is low or the overall connectivity of the scene graph is sparse.
In this experiment, we define three difficulty levels for each evaluation protocol:
  • Image Overlap: overlap ratio per image pair — easy $\geq 50\%$, medium $10\%\leq\cdot<50\%$, hard $10\%$.
  • Scene Sparsity: average pairwise overlap per scene — easy $\geq 10\%$, medium $4\%\leq\cdot<10\%$, hard $4\%$.
Discussion. In Tab. 3, GPT-4o outperforms other VLMs. Both Covis and GPT-4o achieve the most significant improvements over all baselines in hard case scenarios. These hard cases highlight that limited viewpoint overlap and scene sparsity are the primary sources of difficulty.
@@ 6.1 DUSt3R for scene reconstruction
Table 5: Comparison of different graph structures within the DUSt3R framework. We evaluate the Average Pose Error (APE), along with time complexity (in seconds) and memory cost (in GB) per scene. Results are averaged over Gibson test set. The Random graph fails in 3D reconstruction test, making APE not applicable.
@@
Quantitative Analysis of Different Graphs. To evaluate pose estimation and reconstruction quality, we calculate the average pose errors (APE) by matching estimated poses with their closest ground truth poses, as shown in Tab. 5. As expected, using complete and ground truth (GT) graphs yields lower APE, demonstrating more accurate 3D reconstruction. However, obtaining a ground truth graph and using the complete graph incurs high annotation and computational costs. Moreover, random and star graphs fail to produce a reliable 3D map, as shown in Fig. III. Therefore, we find the co-visibility graph to be one of the most suitable options for DUSt3R, balancing memory costs with a reliable 3D map, as shown in Fig. 6. Detailed reconstruction visualizations are available in the supplementary.
@@ 6.2 Automated Training Set Labeling
Figure 7: Reconstructed co-visibility regions during inference. Visual comparison of different graph configurations. Top row: Co-visibility graph for training. Bottom row: High co-visibility graph for training. Columns: Reference image, predicted image, and ground truth target image. In the predicted images, the masked regions are visible as network input, and the brightened areas indicate the reconstructed output. Additionally, the random graph fails in cross-view completion.
@@ 7 Conclusion
Limitation. Co-VisiON is limited to Gibson and HM3D scenes with limited human annotations. Future work will expand and diversify the dataset using additional sources, including real-world environments.
Summary. This paper introduces the Co-Visibility reasONing (Co-VisiON) task, where models predict view connectivity and co-visibility graphs from sparse indoor image sets. We provide a curated dataset with human and automated annotations, show that Co-VisiON challenges existing baselines, and propose Covis - a multi-view model that surpasses pairwise methods by capturing complex spatial relationships. Co-VisiON aims to advance embodied scene understanding research.
@@
This supplementary provides additional details and results that could not fit in the main paper. Specifically, we include: (1) dataset and annotation generation (Gibson and HM3D), (2) top-down visualization of co-visibility graph, (3) evaluation metric details, (4) graph definitions used in DUSt3R and CroCo experiments, (5) ablation studies on Covis, and (6) Sim2Real downstream task details. All experiments and dataset generation were conducted on an HPC cluster with A100 and V100 GPUs, using a single GPU per run.
@@ B.1.2 Human co-visibility annotation
We will release both the automatically labeled and the manually labeled datasets to facilitate future research. The human annotation website will also be released so that anyone can extend the Co-VisiON task to a larger scale and for more diverse scenarios.
@@ G Ablation Study of CoVis
G.1 Zero-shot Evaluation
We further evaluate the generalization ability of our Covis by training on one dataset (either Gibson or HM3D) and testing on both Gibson and HM3D individually. As shown in Tab. I, Multi-view Covis consistently outperforms pairwise Covis in both in-domain and zero-shot settings, demonstrating better robustness to dataset’s domain shifts. We also evaluate the predicted masks using the standard IoU metric (not graph IoU), and achieve an average IoU of 67.3% against the binarized ground-truth masks.
Table I: Zero-shot evaluation of multi-view Covis and pairwise Covis across different training and test sets.